# A Curious Case of Remarkable Resilience to Gradient Attacks via Fully Convolutional and Differentiable Front End with a Skip Connection

### v1

@@ 5.2 Discussion of Results
Although the gradient masking phenomenon is not new, the degree of this masking is quite remarkable for fully differentiable models that do not have gradient-shattering components such as JPEG compression. For example, Tramèr et al. (2020) describe a case when gradient masking was caused by an additional “log-softmaxing” which is known to reduce “gradient flows”. However, in our case, a gradient-masking culprit is a front end, which is a fully-convolutional network with a skip connection and batch-normalization layers (see § A.1 where we describe checking for numerical abnormalities). Both of these architectural components were instrumental to training deep neural networks, which otherwise would suffer from vanishing or exploding gradients (Szegedy et al., 2014; He et al., 2016).
@@
We observed that AutoAttack can be ineffective against randomized ensembles. In particular, on a sample of CIFAR10 images we can achieve an accuracy of 90.8±$\pm 2.5$% (99% CI) for a model that has true robust accuracy under 20%. Is AutoAttack, in principle, weak against randomized ensembles or is there something special about our ensemble models?
First of all, we confirmed that even standard PGD can be extremely effective against randomized ensembles where individual models underwent only standard training using clean data. Indeed, from the Table 1 and Tables 6, 7 in Appendix A.2 we see that accuracy of such ensembles is nearly zero under a standard PGD attack. This was an unexpected finding because such attacks do not use averaging techniques such as EOT and each gradient computation uses a mix of models with different random weights.
@@
2/255 30±$\pm 2$% $\pm 6$% 29±$\pm 5$%
4/255 47±$\pm 2$% 67±$\pm 5$% 42±$\pm 6$%
8/255 52±$\pm 3$% $\pm 6$% 47±$\pm 6$%
@@
A more careful investigation showed that—despite the lack of EOT—because all models in the ensemble share the same architecture, the standard PGD finds adversarial examples that work for all models in the ensemble. To prevent this, we trained ensembles using three different types of backbone models, namely, ResNet-152 (He et al., 2016), wide ResNet-28-10 (Zagoruyko & Komodakis, 2016), and ViT (Dosovitskiy et al., 2021). We did not use any front ends and trained three ensembles using three values of attack $\varepsilon$: $2/255$, $4/255$, and $8/255$. For each model in the ensemble we computed the accuracy under a standard PGD attack (with $\varepsilon=8/255$) and then obtained ensemble averages, which ranged from 30% to 52% depending on the train-time $\varepsilon$.
@@
As expected, the smaller was $\varepsilon$ used during training, the lower was the robust accuracy of the models. However, when we attacked a randomized ensemble using a 50 step PGD with five restarts, we achieved the accuracy close to 65% (see Table 5). Thus, the randomized ensemble of diverse models can fool the standard PGD attack without EOT. In contrast, the accuracy under AutoAttack was slightly lower (as expected) than ensemble-average robust accuracy (see Table 5), thus, confirming that AutoAttack can indeed be very effective against randomized ensembles.
@@ 6 Conclusion
We discovered that training a fully differentiable front end with a frozen classification model can result in the model that is unusually resistant to gradient attacks while retaining accuracy on original data. We provided evidence that this behavior was due to gradient masking. Although the gradient masking phenomenon is not new, the degree of this masking is quite remarkable for fully differentiable models that do not have gradient-shattering components such as JPEG compression or components that are expected to cause diminishing gradients. The objective of this paper is to document this phenomenon and discuss potential consequences for automatic robustness evaluation packages such as AutoAttack (Croce & Hein, 2020a) rather than to establish new SOTA in adversarial robustness.
@@ Appendix A Appendix
A.1 Gradient Tracing: Checking for Numerical Instabilities
@@
To make sure our results were not affected by numerical instabilities (or by exploding and vanishing gradients), we “traced” gradients for ResNet-152 (He et al., 2016) backbone on all three datasets. We used PGD with 50 steps and a single restarted. We did this for all datasets using a sample of 500 images. We compared gradients between two scenarios:
@@
  • Attacking a front-end “ehanced” model that was adversarially trained.
@@
[…]
  • In both cases only a modest fraction (10-30%) of absolute gradient values were small ($<0.01$);
@@
A.2 Additional Experimental Results
Table 6: Accuracy under attack, single model scenario (CIFAR10)
@@ A.2 Additional Experimental Results
Resnet-18 91±$\pm 3$% 0% 1±$\pm 0$% 3±$\pm 2$%
Resnet-18 +front 91±$\pm 3$% 88±$\pm 4$% 75±$\pm 5$% $\pm 1$% 5±$\pm 2$% 0%
Resnet-50 95±$\pm 3$% 0% 0±$\pm 1$% 4±$\pm 3$%
Resnet-50 +front 95±$\pm 3$% 93±$\pm 3$% 78±$\pm 5$% $\pm 2$% 5±$\pm 3$% 0±$\pm 1$%
Resnet-152 94±$\pm 3$% 0% 1±$\pm 2$% 6±$\pm 3$%
Resnet-152 +front 94±$\pm 3$% 92±$\pm 3$% 81±$\pm 4$% $\pm 2$% 7±$\pm 3$% 0%
WideResNet-28-10 96±$\pm 2$% 0% 0% 8±$\pm 3$%
WideResNet-28-10 +front 95±$\pm 3$% 91±$\pm 3$% 65±$\pm 5$% $\pm 1$% 10±$\pm 4$% $\pm 1$%
VIT-16-224 97±$\pm 2$% 0±$\pm 1$% $\pm 1$% 10±$\pm 4$%
VIT-16-224 +front 97±$\pm 2$% 96±$\pm 2$% 85±$\pm 4$% 1±$\pm 2$% 10±$\pm 3$% 0±$\pm 1$%
@@
Resnet-18 3$\times$ 92±$\pm 3$% $\pm 2$% 76±$\pm 5$% $\pm 6$%
Resnet-18 +front 3$\times$ 91±$\pm 4$% 89±$\pm 4$% 70±$\pm 6$% 74±$\pm 5$% 64±$\pm 5$% 3±$\pm 2$%
Resnet-50 3$\times$ 95±$\pm 3$% 3±$\pm 2$% 80±$\pm 4$% 69±$\pm 5$%
Resnet-50 +front 3$\times$ 96±$\pm 2$% 94±$\pm 3$% 71±$\pm 5$% 81±$\pm 4$% 69±$\pm 5$% 5±$\pm 2$%
Resnet-152 3$\times$ 94±$\pm 3$% 3±$\pm 2$% 84±$\pm 4$% 70±$\pm 5$%
Resnet-152 +front 3$\times$ 95±$\pm 2$% 94±$\pm 3$% 71±$\pm 5$% 79±$\pm 5$% 74±$\pm 5$% 6±$\pm 3$%
WideResNet-28-10 3$\times$ 96±$\pm 2$% 3±$\pm 2$% 77±$\pm 5$% 45±$\pm 6$%
WideResNet-28-10 +front 3$\times$ 95±$\pm 3$% 92±$\pm 3$% $\pm 6$% 79±$\pm 4$% $\pm 5$% 3±$\pm 2$%
VIT-16-224 3$\times$ 97±$\pm 2$% $\pm 1$% 78±$\pm 5$% $\pm 6$%
VIT-16-224 +front 3$\times$ 96±$\pm 2$% 96±$\pm 2$% 74±$\pm 5$% 81±$\pm 4$% $\pm 6$% $\pm 1$%
@@
  • We do not run other smooth PGD attacks when the accuracy $\approx 0$% with the standard PGD attack.
Table 7: Detailed Results for CIFAR100
@@
Resnet-18 $\pm 6$% 0±$\pm 1$% $\pm 1$% 4±$\pm 3$%
Resnet-18 +front 68±$\pm 5$% 65±$\pm 5$% 48±$\pm 6$% 2±$\pm 1$% 4±$\pm 2$% 0±$\pm 1$%
Resnet-50 75±$\pm 5$% 1±$\pm 0$% 1±$\pm 2$% 6±$\pm 3$%
Resnet-50 +front 77±$\pm 5$% 70±$\pm 6$% 53±$\pm 6$% $\pm 2$% 7±$\pm 3$% $\pm 1$%
Resnet-152 76±$\pm 5$% 1±$\pm 2$% 2±$\pm 1$% 9±$\pm 3$%
Resnet-152 +front 74±$\pm 5$% 71±$\pm 5$% 51±$\pm 6$% 5±$\pm 2$% 11±$\pm 4$% $\pm 2$%
WideResNet-28-10 76±$\pm 5$% 2±$\pm 1$% 0±$\pm 1$% 7±$\pm 3$%
WideResNet-28-10 +front 75±$\pm 5$% 67±$\pm 5$% 35±$\pm 5$% 2±$\pm 1$% 8±$\pm 3$% 3±$\pm 1$%
VIT-16-224 84±$\pm 5$% $\pm 1$% $\pm 2$% 8±$\pm 3$%
VIT-16-224 +front 84±$\pm 5$% 80±$\pm 5$% 59±$\pm 6$% 3±$\pm 2$% 8±$\pm 4$% $\pm 1$%
@@
Resnet-18 3$\times$ 69±$\pm 5$% 5±$\pm 3$% 51±$\pm 6$% 49±$\pm 5$%
Resnet-18 +front 3$\times$ $\pm 6$% 67±$\pm 5$% 40±$\pm 6$% 54±$\pm 6$% 46±$\pm 6$% 8±$\pm 3$%
Resnet-50 3$\times$ 73±$\pm 5$% 6±$\pm 2$% 57±$\pm 6$% 49±$\pm 6$%
Resnet-50 +front 3$\times$ 75±$\pm 5$% 72±$\pm 6$% 46±$\pm 6$% 58±$\pm 6$% 51±$\pm 6$% 9±$\pm 3$%
Resnet-152 3$\times$ 76±$\pm 5$% 10±$\pm 3$% 58±$\pm 6$% 49±$\pm 6$%
Resnet-152 +front 3$\times$ 75±$\pm 5$% 71±$\pm 5$% 48±$\pm 6$% 57±$\pm 6$% 53±$\pm 6$% 10±$\pm 4$%
WideResNet-28-10 3$\times$ 78±$\pm 5$% 5±$\pm 2$% $\pm 5$% 31±$\pm 5$%
WideResNet-28-10 +front 3$\times$ 76±$\pm 5$% 70±$\pm 5$% 34±$\pm 6$% 52±$\pm 6$% 33±$\pm 6$% 5±$\pm 2$%
VIT-16-224 3$\times$ 85±$\pm 4$% $\pm 2$% $\pm 6$% 46±$\pm 5$%
VIT-16-224 +front 3$\times$ 86±$\pm 4$% 81±$\pm 5$% 54±$\pm 6$% 63±$\pm 5$% 45±$\pm 6$% 3±$\pm 1$%
@@
  • We do not run other smooth PGD attacks when the accuracy $\approx 0$% with the standard PGD attack.
Table 8: Detailed Results for ImageNet
@@
Resnet-18 69±$\pm 9$% 0% $\pm 2$% 23±$\pm 8$%
Resnet-18 +front 70±$\pm 8$% 68±$\pm 9$% 28±$\pm 8$% 2±$\pm 3$% 23±$\pm 8$% 0%
Resnet-50 76±$\pm 8$% 0% 7±$\pm 5$% 40±$\pm 9$%
Resnet-50 +front $\pm 7$% 66±$\pm 9$% 31±$\pm 9$% 8±$\pm 5$% 40±$\pm 10$% 0%
Resnet-152 78±$\pm 8$% 0% 12±$\pm 5$% 46±$\pm 9$%
Resnet-152 +front 78±$\pm 8$% 74±$\pm 9$% 26±$\pm 8$% 16±$\pm 6$% 47±$\pm 9$% 0%
VIT-16-224 83±$\pm 7$% 0% 18±$\pm 8$% 60±$\pm 9$%
VIT-16-224 +front $\pm 8$% 79±$\pm 8$% 48±$\pm 9$% 21±$\pm 8$% 62±$\pm 9$% 0%
@@
Resnet-18 3$\times$ 69±$\pm 9$% 0% $\pm 2$% 23±$\pm 8$%
Resnet-18 +front 3$\times$ 70±$\pm 8$% 68±$\pm 9$% 42±$\pm 10$% 48±$\pm 9$% 33±$\pm 9$% 0±$\pm 2$%
Resnet-50 3$\times$ 76±$\pm 8$% 0% 8±$\pm 4$% 38±$\pm 8$%
Resnet-50 +front 3$\times$ $\pm 7$% 72±$\pm 8$% 50±$\pm 9$% 55±$\pm 9$% 50±$\pm 9$% 0%
Resnet-152 3$\times$ 78±$\pm 8$% 0% 12±$\pm 5$% 44±$\pm 9$%
Resnet-152 +front 3$\times$ $\pm 7$% 74±$\pm 8$% 48±$\pm 10$% 53±$\pm 9$% 53±$\pm 9$% 0%
VIT-16-224 3$\times$ 83±$\pm 7$% 0% 18±$\pm 7$% 58±$\pm 10$%
VIT-16-224 +front 3$\times$ $\pm 8$% 80±$\pm 7$% 63±$\pm 9$% 69±$\pm 9$% 68±$\pm 9$% 0%
@@
  • We do not run other smooth PGD attacks when the accuracy $\approx 0$% with the standard PGD attack.

### v2

@@ 5.2 Discussion of Results
Although gradient masking is a known phenomenon, it is unusual to observe in fully differentiable models without gradient-diminishing or gradient-shattering components. For example, Tramèr et al. (2020) describe a case when gradient masking was caused by an additional “log-softmaxing” which is known to reduce “gradient flows”. However, in our case, a gradient-masking culprit is a front end, which is a fully-convolutional network with a skip connection and batch-normalization layers (see § A.2 where we describe checking for numerical abnormalities). Both of these architectural components were instrumental to training deep neural networks, which otherwise would suffer from vanishing or exploding gradients (Szegedy et al., 2014; He et al., 2016).
Table 5: Calculating the lower bound for the accuracy of the randomized ensemble under a transfer attack with a single source model.
{textenv}
We calculate a lower bound for the expected accuracy of the randomized ensemble under a transfer attack, assuming the following:
  • An attacker creates an attack by using one model at a time, i.e., there is no ensembling of source models;
  • A source model is one of the models used in the randomized ensemble;
  • Individual models in the randomized ensemble are selected with equal probabilities.
According to Table 8 (in § A.1):
  • $A(i\rightarrow i)\approx 0$ (self-attack);
  • $A_{k\rightarrow i}\geq 44.2\%$ if both $i$ and $k$ are different ResNet models (attacking a ResNet using a different ResNet model);
  • $A_{k\rightarrow i}\geq 73.6\%$ if $i$ is a ViT model and $k$ is a ResNet model (attacking a ViT model using a ResNet model);
  • $A_{k\rightarrow i}\geq 65.4\%$ if $i$ is a ResNet model and $k$ is a ViT model (attacking a ResNet model using a ViT model);
  • $A_{k\rightarrow i}\geq 55.9\%$ if both $i$ and $k$ are ViT models (attacking a ViT model using a different ViT model).
First, there is a $10\%$ chance that the source model and the target model are the same. In this case, the average accuracy is near zero. For a source ResNet model, in four cases out of 10, $A_{k\rightarrow i}\geq 44.2\%$ (when a target model is also a ResNet model) and in five cases out of ten $A_{k\rightarrow i}\geq 73.6\%$ (when the target is a ViT). Likewise for a source ViT model, in four cases out of ten $A_{k\rightarrow i}\geq 55.9\%$ and in five cases out of ten $A_{k\rightarrow i}\geq 65.4\%$. Therefore, for a source ResNet model, the lower bound for the expected accuracy is:
$0\%\cdot 0.1+44.2\%\cdot 0.4+73.6\%\cdot 0.5\approx 54.6\%.$
For a source ViT model, the lower bound for the expected accuracy is:
$0\%\cdot 0.1+55.9\%\cdot 0.4+65.4\%\cdot 0.5\approx 55\%$
@@
[…]
First of all, we confirmed that even standard PGD can be extremely effective against randomized ensembles where individual models underwent only standard training using clean data. Indeed, from Table 1 and Tables 10, 11 in Appendix A.3 we see that the accuracy of such ensembles is nearly zero under a standard PGD attack. This was an unexpected finding because such attacks do not use averaging techniques such as EOT and each gradient computation uses a mix of models with different random weights.
@@
A more careful investigation showed that—despite the lack of EOT—because all models in the ensemble share the same architecture, the standard PGD finds adversarial examples that work for all models in the ensemble. To prevent this, we trained ensembles using three different types of backbone models, namely, ResNet-152 (He et al., 2016), wide ResNet-28-10 (Zagoruyko & Komodakis, 2016), and ViT (Dosovitskiy et al., 2021). We did not use any front ends and trained three ensembles using three values of attack $\varepsilon$: $2/255$, $4/255$, and $8/255$. For each model in the ensemble we computed the accuracy under a standard PGD attack (with $\varepsilon=8/255$) and then obtained ensemble averages, which ranged from 30% to 52% depending on the training-time $\varepsilon$.
@@
As expected, the smaller $\varepsilon$ was during training, the lower was the robust accuracy of the models. However, when we attacked a randomized ensemble using a 50-step PGD with five restarts, we achieved the accuracy close to 65% (see Table 6). Thus, the randomized ensemble of diverse models can fool the standard PGD attack without EOT. In contrast, the accuracy under AutoAttack was slightly lower (as expected) than ensemble-average robust accuracy (see Table 6), thus, confirming that AutoAttack can indeed be very effective against randomized ensembles.
Can randomized ensembling offer a practical defense?
While our primary focus is not on designing or benchmarking a specific randomized-ensemble scheme, it is worth discussing whether such an approach could meaningfully boost robustness. To this end, we should first consider plausible attack scenarios. Sitawarin et al. (2024); Qin et al. (2021) showed that randomization thwarts several popular query-based attacks (which we also confirm for the AutoAttack case). Therefore, a typical attacker has to rely on transfer attacks or on hybrid approaches that combine transfer attacks with query-based attacks.
Assume that the defender has $n$ models and makes a prediction by randomly selecting a model according to probabilities $\{p_{i}\}$. The attacker has $m$ models from which they select a source model $k$ to generate an adversarial example. Given such a setup, the expected accuracy of the randomized ensemble is:
$A(k)=\sum_{i=1}^{n}A_{k\rightarrow i}\cdot p_{i},$ (7)
where $A_{k\rightarrow i}$ is the expected accuracy of model $i$ when making a prediction on an adversarial example transferred from model $k$. The best the attacker can do in this situation is to select $k=\operatorname*{arg\,min}_{k=1}^{m}A(k)$ that minimizes the expected accuracy under attack. In the worst case for the defender, the attacker has detailed knowledge about model architectures and parameters and, thus, be able to compute $A(k)$ even without querying the randomized ensemble. Nonetheless, even in the absence of such knowledge, the attacker may be able to estimate $A(k)$ by probing the ensemble for different values of kk—albeit this may not always be practical due to a significant reduction in efficiency (a reliable estimate requires a lot of queries for each $k$).
To make things concrete, let us consider 10 models used in our transferability study for ImageNet (see § A.1 and Table 8). This study includes five ResNet and five ViT models trained using clean data. Because these models were not trained adversarially, even small-radius PGD attacks “drive” their accuracy to zero, i.e., $A(i\rightarrow i)\approx 0$. However, $A(k\rightarrow i)$ is well above zero for $i\neq k$.
To simplify the presentation, we moved the calculations of the expected accuracy under attack to Table 5. In summary, no matter what the source model is, the defender is guaranteed to have the accuracy under attack above 54%. This is about 23 percentage point decrease compared to average accuracy on the clean data, which is 77%. However, it is still markedly better than a nearly 0% accuracy under attack, and importantly, the clean data accuracy is fully retained.
Table 7: Calculating the expected accuracy of the randomized ensemble under the transfer attack that uses the all-model ensemble.
{textenv}
We calculate the expected accuracy of the randomized ensemble under attack, assuming the following:
  • The attacker creates an attack using an ensemble model, which includes all models;
  • Individual models in the randomized ensemble are selected with equal probabilities.
According to Table 8 (in § A.1):
  • $A_{\mbox{ensemble}\rightarrow i}=18.1\%$ if $i$ is a ResNet model;
  • $A_{\mbox{ensemble}\rightarrow i}=26.1\%$ if $i$ is a ViT model.
Because a ViT model is randomly selected 50% of the time (in the ensemble) the expected accuracy is equal to:
$18.1\%\cdot 0.5+26.1\%\cdot 0.5\approx 22\%$
Instead of using individual models, the attacker can use ensembles as source models, which are known to produce stronger attacks (Liu et al., 2017; Gubri et al., 2022). A particularly unfortunate for the defender situation is when the attacker can ensemble all models to generate an attack. In this case (see Table 7 for details), the accuracy under attack is only 22%, which is a $2.5\times$ decrease compared to the case when the attacks are generated using a single model.
It should now be clear that randomization can be a practical defense only if one combines models with low mutual attack transferability. Further exploration of the feasibility of this approach is out of the scope of this work. However, we would like to mention several approaches to achieve this objective. First, Bai et al. (2024) found there is only limited transferability among models trained using adversarial attacks with different strengths, which is also confirmed by our own experiments (see Table 9 in § A.1). Not only transfer is limited when the source model is trained on clean-data, but, surprisingly, adversarial examples coming from more robust models do not work well for less robust models. Second—as discussed in § 2—there are additional training approaches that reduce mutual transferability among models (Sharif et al., 2019; Bui et al., 2021; Sitawarin et al., 2024). Third, one can modify existing architectures in controlled ways, e.g., by changing normalization approaches (Dong et al., 2022) or precision (Fu et al., 2021). This can be used to further improve the architectural diversity of the randomized ensemble.
@@ 6 Conclusion
We discovered that training a fully differentiable front end, which also has a skip connection, with a frozen classification model can result in a model that is unusually resistant to gradient attacks while retaining accuracy on original data. We provided evidence that this behavior is due to gradient masking. Although the gradient masking phenomenon is not new, the degree of this masking is remarkable for fully differentiable models lacking gradient-shattering components like JPEG compression or other elements known to diminish gradients. The objective of this paper is to document this phenomenon and expose weaknesses of AutoAttack (Croce & Hein, 2020a) rather than to establish new SOTA in adversarial robustness. We conclude the paper with a discussion of whether randomized ensembling can function as a practical defense.
7 Acknowledgments
This work was sponsored by DARPA grant HR11002020006.
@@ Appendix A Appendix
A.1 Transfer Attack Study
We have carried out a study of the effectiveness of transfer attacks using two different scenarios, in particular, to assess feasibility of using randomized ensembles as practical defenses. Attacks were generated using a 50-step PGD: $\varepsilon=8/255$ for CIFAR datasets and $\varepsilon=4/255$ for ImageNet. We used the same subsets of CIFAR10, CIFAR100, and ImageNet as in the main experiments.
In the first scenario, we assessed transferability among ten different architectures: Five of the models belong to the ResNet class (from ResNet-18 to ResNet-152) (He et al., 2016) and five to the ViT (Dosovitskiy et al., 2021) class. ViT models differed in their sizes (tiny, small, base) and input patch sizes. In addition to measuring “clean” accuracy and accuracy under a PGD attack, we computed the accuracy under a transfer attack for several source-model architectures. A source model can be either a single model or an ensemble, but a target model is always one of the above-mentioned ResNet or ViT models. The source-models included:
  • The same model as the target model (the case of a self-attack).
  • An individual model whose architecture is different from the target model. To simplify presentation, we only show accuracy for source models producing the strongest attack (i.e., that led to the lowest accuracy). In doing so, we show results for the best ResNet and the best ViT model separately.
  • An ensemble of models. We tested two types of ensembles: (1) a leave-one-out ensemble that does not include the target model, and (2) an all-inclusive ensemble that has all models of a given type (or types). All ensembles were additionally divided into three groups:
    • An ensemble of ResNet models, which has either all ResNet models (an inclusive ensemble) or all ResNet models except the target model (a leave-one-out ensemble);
    • An ensemble of ViT models, which has either all ViT models (an inclusive ensemble) or all ViT models except the target model (a leave-one-out ensemble);
    • An all-models ensemble, which has either all 10 models (an inclusive ensemble) or all models except the target one (a leave-one-out ensemble);
To simplify presentation, we only show summary results obtained by averaging accuracies for all target models of the same class. From Table 8, we can see that the accuracy under the self-attack is always under 5%. This is expected because models were not trained adversarially. However, attacks from different architectures are less successful, especially attacks from ResNet to ViT models and vice versa. Moreover, generating attacks using ensembles produces stronger attacks, which is well known from the literature (Liu et al., 2017; Gubri et al., 2022). Finally, we note that on ImageNet we observe the lowest transferability of attacks.
Table 8: Transfer attack accuracy among different model architectures (see table notes). We show averages for ResNet and ViT models.
CIFAR10 CIFAR100 ImageNet
source target ResNet ViT ResNet ViT ResNet ViT
Source model and target models are different
No attack 92.7% 96.5% 73.0% 80.3% 74.7% 79.1%
Self-attack 0.3% (-92%) 0.6% (-96%) 2.2% (-71%) 1.4% (-79%) 0.0% (-75%) 3.1% (-76%)
Best ResNet 33.2% (-60%) 80.3% (-16%) 31.5% (-41%) 59.9% (-20%) 44.2% (-31%) 73.6% (-5%)
Best ViT 67.0% (-26%) 12.9% (-84%) 50.4% (-23%) 16.9% (-63%) 65.4% (-9%) 55.9% (-23%)
Leave-One-Out (LOT) ensembles: target models are excluded
ResNet ensemble 24.3% (-68%) 63.8% (-33%) 19.6% (-53%) 46.6% (-34%) 32.9% (-42%) 71.1% (-8%)
ViT ensemble 53.9% (-39%) 10.0% (-87%) 39.2% (-34%) 12.3% (-68%) 60.6% (-14%) 55.2% (-24%)
All ensemble 25.6% (-67%) 39.0% (-58%) 20.9% (-52%) 23.2% (-57%) 34.9% (-40%) 59.1% (-20%)
All-inclusive ensembles: All models of a given type are included
ResNet ensemble 15.4% (-77%) 62.8% (-34%) 11.4% (-62%) 46.9% (-33%) 13.8% (-61%) 71.0% (-8%)
ViT ensemble 52.9% (-40%) 6.4% (-90%) 38.8% (-34%) 6.8% (-73%) 60.5% (-14%) 15.4% (-64%)
All ensemble 17.8% (-75%) 35.3% (-61%) 13.9% (-59%) 17.8% (-63%) 18.1% (-57%) 26.1% (-53%)
  • We experiment with five ResNet models and five ViT models: Best ResNet and Best ViT denote the most effective attack using ResNet or ViT.
  • Model are trained using “clean” data, i.e., without adversarial attacks.
  • In round brackets we show the loss of accuracy compared to “clean” accuracy (no attack), which we round to a nearest integer.
In our next experiment, we tested five ResNet and two ViT architectures on CIFAR10. Each architecture variant was trained using a different scenario. Scenarios include (1) training on clean data and (2) adversarial training with different attack strengths: from $\varepsilon=2/255$ to $8/255$. The attack radius was $\varepsilon=8/255$.
From Table 9, we can see that transferring an attack from a model trained on clean data to a model trained with the attack radius $\varepsilon=2/255$ reduces the accuracy by only 6.7 percentage points. This is totally unsurprising since the clean-data model is not supposed to produce strong attacks for robust models. However, transferring attacks in the other direction—i.e., from the model trained adversarially with $\varepsilon=2/255$ to a model trained on clean data—yields 34.7% accuracy, still massively higher than the 0.4% achieved under the self-attack.
Table 9: CIFAR10: average transfer attack accuracy among models with the same architecture, but trained with different attack strengths: In round brackets we show the loss of accuracy compared to “clean” accuracy (no attack). Diagonal row shows the self-attack accuracy (attack is generated using the same model). We tested five ResNet and two VIT architectures.
$\varepsilon$-train No attack Accuracy under transfer attack
source target no adv. train. 2/255 4/255 8/255
no adv. train 93.7% 0.4% (-93.4%) 83.1% (-6.7%) 83.1% (-3.8%) 75.3% (-2.6%)
2/255 89.8% 34.7% (-59.0%) 34.6% (-55.3%) 60.9% (-26.0%) 65.9% (-12.0%)
4/255 86.9% 52.2% (-41.5%) 51.8% (-38.0%) 46.9% (-40.0%) 60.4% (-17.5%)
8/255 77.9% 72.5% (-21.2%) 67.1% (-22.7%) 61.7% (-25.2%) 51.8% (-26.1%)
A.2 Gradient Tracing: Checking for Numerical Instabilities
@@
To make sure our results were not affected by numerical instabilities (or by exploding and vanishing gradients), we “traced” gradients for ResNet-152 (He et al., 2016) backbone on all three datasets. We used PGD with 50 steps and a single restart. We did this for all datasets using a sample of 500 images. We compared gradients between two scenarios:
@@
  • Attacking a front-end “enhanced” model that was adversarially trained.
@@
  • No gradient was ever NaN or $\pm\infty$;
  • In both cases only a modest fraction (10-30%) of absolute gradient values were small (<0.01<0.01);
@@
A.3 Additional Experimental Results
Table 10: Detailed results for CIFAR10 (see table notes).
@@ A.3 Additional Experimental Results
[…]
@@
[…]
@@
  • We do not run additional PGD attacks (except 0-order PGD) when the accuracy $\approx 0$% with the standard PGD attack.
Table 11: Detailed results for CIFAR100 (see table notes).
@@
[…]
@@
[…]
@@
  • We do not run additional PGD attacks (except 0-order PGD) when the accuracy $\approx 0$% with the standard PGD attack.
Table 12: Detailed results for ImageNet (see table notes).
@@
[…]
@@
[…]
@@
  • We do not run additional PGD attacks (except 0-order PGD) when the accuracy $\approx 0$% with the standard PGD attack.